OpenStudy (anonymous):

Rewrite the expression as a trigonometric function of a single angle measure. (tan4θ + tan2θ)/(1 - tan4θ tan2θ)

OpenStudy (anonymous):

\[ \tan ( x + y)= \frac { \tan x + \tan y }{1-\tan x \tan y} \] Apply this formula.

OpenStudy (anonymous):

The answer should be \[ \tan ( 4 \theta + 2 \theta)=\tan(6 \theta) \]

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

yes, i'm looking over it. thank you!

OpenStudy (anonymous):

yw